What would be the difference in Average Flowtime if you did the following jobs by Shortest Processing Time instead of how they are listed?

Job   Processing Time   Days Due from Today
A   16   22
B   21   30
C   22   36
D   14   28
E   10   18
Question 3 options:

Average flowtime would decrease by 6.4


Average flowtime would decrease by 10.6


Average flowtime would decrease by 10.0


Average flowtime would decrease by 13.8


Average flowtime would decrease by 12.6
Answer #1

Average flowtime would decrease by 10.0

At present the overall flow time of the Jobs as per sequence would be

A - 16

B - 37

C - 59

D - 73

E - 83

Total flow time = (16+37+69+73+83)

= 268 Days

Average flow time = 268/5

= 53.6

In case of shortest processing time, the sequence of processing would be

E-D-A-B-C

The overall flow time of the Jobs as per sequence

E - 10

D - 24

A - 40

B - 61

C - 83

Total flow time = (10+24+40+61+83)

= 218

Average flow time = 218/5

= 43.6

Total Change = 53.6-43.6

= 10